What is the equilibrium Pb2+ concentration when 2.61 L of a 0.103 M lead acetate solution are mixed with 2.82 L of a 0.370 M ammonium bromidesolution?

[Pb2+] =  M

The solubility of AgCl in water is  1.34×10-5 M, but increases in the presence of thiosulfate ion. Calculate the solubility of AgCl when the solid is added to a solution that is 0.520 M in thiosulfate ion. For AgCl, Ksp = 1.80×10-10 and for Ag(S2O3)23-, Kf = 2.00×1013.

Solubility =  M
Answer #1

we know, Ksp of lead boumide, PbBr2 = 6:60x10° (837 = (0.340x2:82) C 0.103) 52:61) 2.61+ 2.822 11043 -0.268 5143 -0.142 m 3.1 Kb = (Ag (8303)2]* cc) 3600 22 00:520-23) 60 = 0 01520-27 31.2-120X = ? 121X=3112 X=0257M The solu bllity. Of Age = 012571.